Hello,

I need help solving this problem using the formula h=vt-16t^2:

An object is fired into the air with a velocity of 72 feet per second. How many seconds will it take for the object to hit the ground?

Thank you


Answer by scott8148(3749)

the object hits the ground when h=0 ... this is also where the object starts; so t=0 is a solution

0=72t-16t^2 ... factoring gives 0=16t(4.5-t) ... 0=16t and 0=4.5-t ... t=4.5